# UniRAG: Universal Retrieval Augmentation for Multi-Modal Large Language Models

### v2

@@ Abstract
Recently, Multi-Modal (MM) Large Language Models (LLMs) have unlocked many complex use-cases that require MM understanding (e.g., image captioning or visual question answering) and MM generation (e.g., text-guided image generation or editing) capabilities. To further improve the output fidelity of MM-LLMs we introduce UniRAG, a plug-and-play technique that adds relevant retrieved information to prompts as few-shot examples during inference. Unlike the common belief that Retrieval Augmentation (RA) mainly improves generation or understanding of uncommon entities, our evaluation results on the MSCOCO dataset with common entities show that both proprietary models like GPT-4o and Gemini-Pro and smaller open-source models like LLaVA, LaVIT, and Emu2 significantly enhance their generation quality when their input prompts are augmented with relevant information retrieved by MM retrievers like UniIR models. All the code required for reproducing our results will be released upon acceptance.
@@
for Multi-Modal Large Language Models
@@
Figure 1: An Overview of the UniRAG technique with image captioning (blue) and image generation (green) tasks. UniRAG retrieves relevant image-text pairs and adds them as few-shot examples to the MM-LLM’s input prompt.
@@ 1 Introduction
Recent advancements in Multi-Modal Large Language Models (MM-LLMs), encompassing both proprietary models like GPT-4o OpenAI et al. (2024), Gemini-Pro Team et al. (2024), DALL-E Ramesh et al. (2021) and Parti Yu et al. (2022) and open-source models such as LLaVA Liu et al. (2023c, b), LaVIT Jin et al. (2023) and Emu2 Sun et al. (2023), are instrumental in bridging the gap between different modalities. These models have demonstrated remarkable human-like efficacy and achieved state-of-the-art effectiveness in various benchmark assessments Li and Lu (2024).
@@
However, like LLMs, MM-LLMs often struggle to produce accurate results on lesser-known or recent topics Huang et al. (2024); Bai et al. (2024). This limitation occurs because models generate responses solely based on their training data, lacking access to external information during inference. Given the rapidly changing nature of the world, it is unrealistic for models to correctly address every query using only pre-trained knowledge.
@@
As a workaround, techniques like Retrieval Augmented Generation (RAG) have gained significant attention in recent years Gao et al. (2023). In-context RAG Ram et al. (2023) particularly relies on the learning capabilities of models to incorporate the latest knowledge and guide them in generating more relevant results during inference. Recent studies have applied RAG to multi-modal tasks, including visual question answering, text-guided image generation and editing, and image captioning Yasunaga et al. (2022); Lin and Byrne (2022); Chen et al. (2022a); Yu et al. (2023); Hu et al. (2023). However, these approaches mainly focus on training models with large image-text datasets, leaving the potential for plug-and-play in-context RAG in MM applications largely unexplored. As black-box MM-LLMs become more prevalent, addressing this gap is essential for their reliable adoption in applications that cannot tolerate hallucinations.
@@
This paper introduces the plug-and-play UniRAG technique, designed to enhance the fidelity of model outputs. UniRAG integrates RA with MM-LLMs to guide the generation process using relevant in-context information. By incorporating interleaved image-text pairs as few-shot examples during inference, UniRAG serves as a model-agnostic approach that effectively teaches out-of-the-box MM-LLMs about various task intents and uncommon entities. Our evaluation in this work primarily focuses on image captioning (image-to-text) and image generation (text-to-image) tasks.
@@
UniRAG adopts a two-stage retrieval and generation approach. For retrieval we use the UniIR Wei et al. (2023) retriever, which is trained with diverse MM datasets to retrieve heterogeneous outputs in both text and image modalities. We mainly adopt UniIR’s CLIP Score Fusion and BLIP Feature Fusion models which are instruction-tuned using CLIP Radford et al. (2021) and BLIP-2 Li et al. (2022), respectively. In the generation stage, we use a variety of off-the-shelf proprietary and open-source MM-LLMs models. In particular, we employ LLaVA, GPT-4o and Gemini-Pro for image captioning and LaVIT and Emu2 for image generation. We guide these MM-LLMs using zero-shot and few-shot prompting techniques to demonstrate their enhanced generation capabilities due to access to relevant examples during inference.
@@
For our evaluations, we mainly utilize the MSCOCO Chen et al. (2015) dataset from the M-BEIR Wei et al. (2023) benchmark. We assess caption and image generation using M-BEIR’s image-to-text and text-to-image tasks, respectively. By combining RA with MM-LLMs, we achieve significant improvements over baseline effectiveness: an average increase of about  9 percentage points in SPICE Anderson et al. (2016) for image captioning and a reduction of 25 units in Fréchet Inception Distance (FID) Heusel et al. (2017) for image generation. Additional evaluations using M-BEIR’s Fashion200k dataset demonstrate that UniRAG is even more effective in domain-specific applications. To summarize, our main contributions in this paper are as follows:
@@
  • Introducing the plug-and-play UniRAG technique, which combines MM retrievers with MM-LLMs using in-context RAG.
@@
  • Extensive evaluation of UniRAG’s effectiveness for image captioning and image generation tasks using five MM-LLMs including both open-source and proprietary models.
@@ 2 Related Work
Multi-Modal Models and Retrieval Augmentation (Vision-Language):
Recently, a number of models have been developed that integrate vision and language understanding Ramesh et al. (2021); Radford et al. (2021); Yu et al. (2022); Li et al. (2022); Liu et al. (2023c, b); Jin et al. (2023); Sun et al. (2023); Zhai et al. (2023); OpenAI et al. (2024); Team et al. (2024). These models typically use modality-specific encoders trained on large image-text datasets to form a unified feature space. More details on the MM-LLMs selected for evaluation can be found in Section 3.2.
Similar to traditional LLMs, MM-LLMs also struggle with generating content about lesser-known entities or uncommon combinations of common ones Chen et al. (2022b); Blattmann et al. (2022). By leveraging retrieval, MM-LLMs can be guided to produce more accurate and high-fidelity outputs. UniIR Wei et al. (2023) introduces MM retriever models that are instruction-tuned with diverse MM datasets to perform heterogeneous retrieval tasks involving both text and images. RA-CM3 Yasunaga et al. (2022), MuRAG Chen et al. (2022a), CM3Leon Yu et al. (2023) and REVEAL Hu et al. (2023) employ a retriever-generator workflow for RAG in various MM tasks. However, they all require training generator and/or retriever models on extensive MM datasets.
In contrast, UniRAG incorporates an MM retriever during inference, facilitating in-context RAG for black-box LLMs. The work by Liu et al. (2023a), which presents an RA chain-of-thought method, is closest to ours. However, unlike UniRAG, it focuses on improving the effectiveness of MM-LLMs for visual question answering.
@@ 3.1 Retrieval
The retrieval stage is where for a given query its top $k$ most relevant candidates are extracted from an MM database. In this stage query and candidate modalities can be any combination of $\{text,image\}\rightarrow\{text,image\}$. Specifically, for an image-to-text task such as caption generation, the query is an image, and retrieved candidates are captions; while for a text-to-image task (e.g., image generation), the query is a caption and retrieved candidates are images.
To use image-text pairs as in-context examples in the next stage, we need to convert the retrieved single-modality candidates into pairs. This involves treating each retrieved candidate as a query to find its complementary candidate based on two criteria: a) it must have the opposite modality (text for an image candidate and vice versa) and b) it must differ from the original query to prevent revealing the ground-truth answer. The retriever models used for our experiments are further explained in Section 4
@@ 3.2 Generation
The generation stage is where the MM-LLM generates the required output in zero-shot Kojima et al. (2023) or few-shot Brown et al. (2020) settings. Similar to the retrieval stage, in the zero-shot setting (baseline 1), the query is only present in the input prompt; while in the few-shot setting, in-context examples are additionally included in the prompt. In this setting, examples are obtained using one of the following methods: 1) random selection of queries and their ground truth candidates from the dataset (baseline 2); 2) in-context RAG with retrieved candidate pairs from the previous stage. We leverage various MM-LLMs for caption and image generation tasks, as detailed in Section 4.
@@ 4 Experimental Setup
Zeroshot CLIP-SF BLIP-FF Random
@@ 4.1 Selected Models
We use UniIR’s CLIP Score Fusion (CLIP-SF) and BLIP Feature Fusion (BLIP-FF) models as MM retrievers. Experimental results from the original paper indicate that these instruction-tuned UniIR models significantly outperform their pre-trained CLIP and BLIP2 baselines across various tasks and configurations, including image and caption retrieval tasks Wei et al. (2023).
@@
LLaVA is an open-source MM model based on Vicuna Chiang et al. (2023), utilizing CLIP as its visual encoder. It is fine-tuned on MM instruction-following data generated by GPT-4 Liu et al. (2023c). LLaVA’s ability to understand and follow human intent makes it well-suited for RAG caption generation. Figures 25 in Appendix A show detailed prompts used for caption generation.
Unlike image captioning, which only requires MM understanding, image generation necessitates models with both MM understanding and generation capabilities. To meet this need, we selected LaVIT and Emu2-Gen from a survey of MM-LLMs Zhang et al. (2024).
LaVIT is an open-source model based on LLaMA-7B Touvron et al. (2023) that employs a visual tokenizer to convert images into discrete visual tokens. This allows it to process both text and image inputs using a unified objective for next image/text token prediction. The authors report a Frechet Inception Distance (FID) of 7.4 Heusel et al. (2017) on the MSCOCO dataset, indicating its strong performance in image generation tasks.
Emu2-Gen is another open-source model, fine-tuned from Emu2 for image generation. Like LaVIT, it uses next token prediction across all modalities and incorporates a variant of CLIP as its visual encoder along with LLaMA-33B as its backbone. It can accept interleaved inputs of images, text, and video, and features a visual decoder for generating images. While its zero-shot performance on various MM tasks is lower than that of Emu Dai et al. (2023), Flamingo Alayrac et al. (2022), and others, it excels in few-shot settings, making it suitable for image generation with RAG.
Table 2 displays visual examples of image generation using LaVIT and Emu2-Gen in zero-shot, random few-shot, and RAG few-shot settings (using CLIP-SF as the retriever).
@@
Zeroshot CLIP-SF BLIP-FF Random
@@
4.2 Dataset
We evaluate UniRAG on image-to-text and text-to-image tasks from the M-BEIR Wei et al. (2023) benchmark using its MSCOCO and Fashion200k datasets. To ensure generic retrieval, we retrieve from M-BEIR’s global candidate pool, which contains over 5.5 million candidates from 10 different datasets, instead of using dataset-specific pools.
The MSCOCO test set includes about 25k captions for 5k unique images of common objects. For caption generation, we use all 5k images as queries. However, due to the lengthy image generation time (over 12 seconds per image), we randomly sample one caption query for each image. To maintain consistency, all image generation runs use this same sample set. We analyze the effects of this sampling in detail in section 5.2.1. To demonstrate UniRAG’s effectiveness in domain-specific tasks, we also evaluate it on the test set of Fashion200k dataset from M-BEIR. This set includes about 1.7k caption and 4.9k image queries, all focused on the fashion domain.
@@ 4.3 Configuration Details
We use LLaVA’s default configuration  [^1] https://github.com/huggingface/transformers/blob/main/src/transformers/models/llava/configuration_llava.py for all its inferences, with a batch size of four for zero-shot and two for few-shot generation. Zero-shot caption generation on a single NVIDIA RTX A6000 GPU averages around 5 seconds, while few-shot takes 2-3 times longer. For Gemini-Pro and GPT-4o, we utilize Vertex AI’s generate content and OpenAI’s chat completion APIs, respectively, setting their max_new_token parameter to 400. More details on their cost estimates can be found in Appendix B. For LaVIT, we follow the sample configuration provided  [^2] https://github.com/jy0205/LaVIT/blob/main/LaVIT/text2image_synthesis.ipynb, generating a 1024×1024 image in about 12 seconds on a single NVIDIA RTX A6000 GPU. Similarly, we use Emu2-Gen’s default settings  [^3] https://github.com/baaivision/Emu/tree/main/Emu2#emu2-gen. Since it has 37 billion parameters, we run its "bf16" variant on an NVIDIA H100 SXM from the Lambda GPU provider. Image generation with this setting takes about 8 seconds for zero-shot and one-shot, and 13 seconds for five-shot. We encountered CUDA out-of-memory errors for 36 out of 5k captions during five-shot image generation with Emu2-Gen, so we limited the prompt to the first four pairs in those cases.
@@
Table 3: Caption generation with LLaVA, Gemini-Pro and GPT-4o models on the MSCOCO’s dataset using UniIR’s CLIP-SF and BLIP-FF as retrievers. Column $k\in\{0,1,5\}$ shows the number of few-shot examples in each run.

### v3

# UniRAG: Universal Retrieval Augmentation for Large Vision Language Models

@@ Abstract
Recently, Large Vision Language Models (LVLMs) have unlocked many complex use cases that require Multi-Modal (MM) understanding (e.g., image captioning or visual question answering) and MM generation (e.g., text-guided image generation or editing) capabilities. To further improve the output fidelity of LVLMs we introduce UniRAG, a plug-and-play technique that adds relevant retrieved information to prompts as few-shot examples during inference. Unlike the common belief that Retrieval Augmentation (RA) mainly improves generation or understanding of uncommon entities, our evaluation results on the MSCOCO dataset with common entities show that both proprietary models like GPT-4o and Gemini-Pro and smaller open-source models like LLaVA, LaVIT, and Emu2 significantly enhance their generation quality when their input prompts are augmented with relevant information retrieved by Vision-Language (VL) retrievers like UniIR models. All the necessary code to reproduce our results is available at https://github.com/castorini/UniRAG
@@
for Large Vision Language Models
@@
Figure 1: An Overview of the UniRAG technique with image captioning (blue) and image generation (green) tasks. UniRAG retrieves relevant image-text pairs and adds them as few-shot examples to the LVLM’s input prompt.
@@ 1 Introduction
Recent advancements in Large Vision Language Models (LVLMs), encompassing both proprietary models like GPT-4o OpenAI et al. (2024), Gemini-Pro Team et al. (2024), DALL-E Ramesh et al. (2021) and Parti Yu et al. (2022) and open-source models such as LLaVA Liu et al. (2023c, b), LaVIT Jin et al. (2023) and Emu2 Sun et al. (2023), are instrumental in bridging the gap between different modalities. These models have demonstrated remarkable human-like efficacy and achieved state-of-the-art effectiveness in various benchmark assessments Li and Lu (2024).
@@
However, like LLMs, VLMs often struggle to produce accurate results on lesser-known or recent topics Huang et al. (2024); Bai et al. (2024). This limitation occurs because models generate responses solely based on their training data, lacking access to external information during inference. Given the rapidly changing nature of the world, it is unrealistic for models to correctly address every query using only pre-trained knowledge.
@@
As a workaround, techniques like Retrieval Augmented Generation (RAG) have gained significant attention in recent years Gao et al. (2023). In-context RAG Ram et al. (2023) particularly relies on the learning capabilities of models to incorporate the latest knowledge and guide them in generating more relevant results during inference. Recent studies have applied RAG to VL tasks, including Visual Question Answering (VQA), text-guided image generation and editing, and image captioning Yasunaga et al. (2022); Lin and Byrne (2022); Chen et al. (2022a); Yu et al. (2023); Hu et al. (2023). However, these approaches mainly focus on training models with large image-text datasets, leaving the potential for plug-and-play in-context RAG in VL applications largely unexplored. As black-box LVLMs become more prevalent, addressing this gap is essential for their reliable adoption in applications that cannot tolerate hallucinations.
@@
This paper introduces the plug-and-play UniRAG technique, designed to enhance the fidelity of model outputs. UniRAG integrates RA with LVLMs to guide the generation process using relevant in-context information. By incorporating interleaved image-text pairs as few-shot examples during inference, UniRAG serves as a model-agnostic approach that effectively teaches out-of-the-box LVLMs about various task intents and uncommon entities. Our evaluation in this work primarily focuses on image captioning (image-to-text) and image generation (text-to-image) tasks.
@@
UniRAG adopts a two-stage retrieval and generation approach. For retrieval we use the UniIR Wei et al. (2023) retriever, which is trained with diverse VL datasets to retrieve heterogeneous outputs in both text and image modalities. We mainly adopt UniIR’s CLIP Score Fusion and BLIP Feature Fusion models which are instruction-tuned using CLIP Radford et al. (2021) and BLIP-2 Li et al. (2022), respectively. In the generation stage, we use a variety of off-the-shelf proprietary and open-source LVLM models. In particular, we employ LLaVA, GPT-4o and Gemini-Pro for image captioning and LaVIT and Emu2 for image generation. We guide these LVLMs using zero-shot and few-shot prompting techniques to demonstrate their enhanced generation capabilities due to access to relevant examples during inference.
@@
For our evaluations, we mainly utilize the MSCOCO Chen et al. (2015) dataset from the M-BEIR Wei et al. (2023) benchmark. We assess caption and image generation using M-BEIR’s image-to-text and text-to-image tasks, respectively. By combining RA with LVLMs, we achieve significant improvements over baseline effectiveness: an average increase of about  9 percentage points in SPICE Anderson et al. (2016) for image captioning and a reduction of 25 units in Fréchet Inception Distance (FID) Heusel et al. (2017) for image generation. Additional evaluations using M-BEIR’s Fashion200k dataset demonstrate that UniRAG is even more effective in domain-specific applications. To summarize, our main contributions in this paper are as follows:
@@
  • Introducing the plug-and-play UniRAG technique, which combines VL retrievers with LVLMs using in-context RAG.
@@
  • Assessing the effectiveness of UniRAG in image captioning and generation tasks using five LVLMs, including both open-source and proprietary models.
@@ 2 Related Work
Vision-Language Models and Retrieval Augmentation:
Recently, a number of models have been developed that integrate vision and language understanding Ramesh et al. (2021); Radford et al. (2021); Yu et al. (2022); Li et al. (2022); Liu et al. (2023c, b); Jin et al. (2023); Sun et al. (2023); Zhai et al. (2023); OpenAI et al. (2024); Team et al. (2024). These models typically use modality-specific encoders trained on large image-text datasets to form a unified feature space. More details on the LVLMs selected for evaluation can be found in Section 3.2.
Similar to traditional LLMs, LVLMs also struggle with generating content about lesser-known entities or uncommon combinations of common ones Chen et al. (2022b); Blattmann et al. (2022). By leveraging retrieval, LVLMs can be guided to produce more accurate and high-fidelity outputs. UniIR Wei et al. (2023) introduces VL retriever models that are instruction-tuned with diverse VL datasets to perform heterogeneous retrieval tasks involving both text and images. RA-CM3 Yasunaga et al. (2022), MuRAG Chen et al. (2022a), CM3Leon Yu et al. (2023) and REVEAL Hu et al. (2023) employ a retriever-generator workflow for RAG in various VL tasks. However, they all require training generator and/or retriever models on extensive VL datasets.
In contrast, UniRAG incorporates a VL retriever during inference, facilitating in-context RAG for black-box LLMs. The work by Liu et al. (2023a) is the closest to ours, introducing an RA Chain-of-Thought (CoT) method. However, unlike UniRAG, it focuses on enhancing LVLMs for VQA. Their approach selects the most relevant task from each VQA benchmark’s training/test set as a few-shot CoT demonstration. By comparison, UniRAG does not rely on task-specific examples to guide the generator. Instead, it retrieves relevant information from an external database to enrich the model’s knowledge during inference. This task-agnostic design allows UniRAG to benefit various image-text tasks, including VQA, image captioning, image generation, and Optical Character Recognition (OCR), by incorporating semantically relevant image-text pairs as few-shot examples.
@@ 3.1 Retrieval
The retrieval stage is where for a given query its top $k$ most relevant candidates are extracted from a VL database. In this stage query and candidate modalities can be any combination of $\{text,image\}\rightarrow\{text,image\}$. Specifically, for an image-to-text task such as caption generation, the query is an image, and retrieved candidates are captions; while for a text-to-image task (e.g., image generation), the query is a caption and retrieved candidates are images.
To use image-text pairs as in-context examples in the next stage, we need to convert the retrieved single-modality candidates into pairs. This involves treating each retrieved candidate as a query to find its complementary candidate based on two criteria: a) it must have the opposite modality (text for an image candidate and vice versa) and b) it must differ from the original query to prevent revealing the ground-truth answer. The retriever models used for our experiments are further explained in Section 4.
@@ 3.2 Generation
The generation stage is where the LVLM generates the required output in zero-shot Kojima et al. (2023) or few-shot Brown et al. (2020) settings. Similar to the retrieval stage, in the zero-shot setting (baseline 1), the query is only present in the input prompt; while in the few-shot setting, in-context examples are additionally included in the prompt. In this setting, examples are obtained using one of the following methods: 1) random selection of queries and their ground-truth candidates from the dataset (baseline 2); 2) in-context RAG with retrieved candidate pairs from the previous stage. We leverage various LVLMs for caption and image generation tasks, as detailed in Section 4.
@@ 4 Experimental Setup
Zero-shot CLIP-SF BLIP-FF Random
@@ 4.1 Selected Models
We use UniIR’s CLIP Score Fusion (CLIP-SF) and BLIP Feature Fusion (BLIP-FF) models as VL retrievers. Experimental results from the original paper indicate that these instruction-tuned UniIR models significantly outperform their pre-trained CLIP and BLIP2 baselines across various tasks and configurations, including image and caption retrieval tasks Wei et al. (2023).
@@
LLaVA is an open-source LVLM based on Vicuna Chiang et al. (2023), utilizing CLIP as its visual encoder. It is fine-tuned on VL instruction-following data generated by GPT-4 Liu et al. (2023c). LLaVA’s ability to understand and follow human intent makes it well-suited for RAG caption generation. Figures 2, 3, 4 and 5 in Appendix A show detailed prompts used for caption generation.
Unlike image captioning, which only requires MM understanding, image generation necessitates models with both MM understanding and generation capabilities. To meet this need, we selected LaVIT and Emu2-Gen from a survey of LVLMs Zhang et al. (2024).
LaVIT is an open-source model based on LLaMA-7B Touvron et al. (2023) that employs a visual tokenizer to convert images into discrete visual tokens. This allows it to process both text and image inputs using a unified objective for next image/text token prediction. The authors report a Frechet Inception Distance (FID) of 7.4 Heusel et al. (2017) on the MSCOCO dataset, indicating its strong capability in image generation tasks.
Emu2-Gen is another open-source model, fine-tuned from Emu2 for image generation. Like LaVIT, it uses next token prediction across all modalities and incorporates a variant of CLIP as its visual encoder along with LLaMA-33B as its backbone. It can accept interleaved inputs of images, texts, and videos, and features a visual decoder for generating images. While its zero-shot effectiveness on various VL tasks is lower than that of Emu Dai et al. (2023), Flamingo Alayrac et al. (2022), and others, it excels in few-shot settings, making it suitable for image generation with RAG.
Table 2 displays visual examples of image generation using LaVIT and Emu2-Gen in zero-shot, random few-shot, and RAG few-shot settings (using UniIR retrievers).
@@
Zero-shot CLIP-SF BLIP-FF Random
@@
4.2 Datasets
We evaluate UniRAG on image-to-text and text-to-image tasks using the MSCOCO and Fashion200k datasets from the M-BEIR Wei et al. (2023) benchmark. To ensure broad retrieval, we use M-BEIR’s global candidate pool of over 5.5 million entries, including images and texts from all 10 datasets. This approach allows us to evaluate in a more realistic scenario with a diverse and extensive external database, rather than limiting retrieval to each dataset’s local corpus.
The MSCOCO test set includes about 25k captions for 5k unique images of common objects. For caption generation, we use all 5k images as queries. However, due to the lengthy image generation time (over 12 seconds per image), we randomly sample one caption query for each image. To maintain consistency, all image generation runs use this same sample set. We analyze the effects of this sampling in detail in Section 5.3. To demonstrate UniRAG’s effectiveness in domain-specific tasks, we also evaluate it on the test set of Fashion200k dataset from M-BEIR. This set includes about 1.7k caption and 4.9k image queries, all from the fashion domain.
@@ 4.3 Configuration Details
Table 3: Caption generation with LLaVA, Gemini-Pro and GPT-4o models on the MSCOCO dataset using UniIR’s CLIP-SF and BLIP-FF as retrievers. Column $k\in\{0,1,5\}$ shows the number of few-shot examples in each run.
We use LLaVA’s default configuration [^1] https://github.com/huggingface/transformers/blob/main/src/transformers/models/llava/configuration_llava.py for all its inferences, with a batch size of four for zero-shot and two for few-shot generation. Zero-shot caption generation on a single NVIDIA RTX A6000 GPU averages around 5 seconds, while few-shot takes 2-3 times longer. For Gemini-Pro and GPT-4o, we utilize Vertex AI’s generate content and OpenAI’s chat completion APIs, respectively, setting their max_new_token parameter to 400. More details on their cost estimates can be found in Appendix B. For LaVIT, we follow the sample configuration provided, [^2] https://github.com/jy0205/LaVIT/blob/main/LaVIT/text2image_synthesis.ipynb generating a 1024×1024 image in about 12 seconds on a single NVIDIA RTX A6000 GPU. Similarly, we use Emu2-Gen’s default settings. [^3] https://github.com/baaivision/Emu/tree/main/Emu2#emu2-gen Since it has 37 billion parameters, we run its “bf16” variant on an NVIDIA H100 SXM from the Lambda GPU provider. Image generation with this setting takes about 8 seconds for zero-shot and one-shot, and 13 seconds for five-shot. We encountered CUDA out-of-memory errors for 36 out of 5k captions during five-shot image generation with Emu2-Gen, so we limited the prompt to the first four pairs in those cases.
5 Evaluation Results and Analysis